# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v4

@@ Appendix C Proof of the approximate proposition
First, we derive the expected bound for each independent element $y_{t}^{(i)}$ in any candidate forecast $\hat{Y}_{f}^{(i)}=\{y_{1}^{(i)},y^{(i)},\cdots,y_{c}^{(i)}\}$.
Since $y_{t}^{(i)}\sim\hat{p}(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})$, we know from the (A2) that $\{y_{t}^{(i)}\}_{i=1}^{c}$ are i.i.d. samples from the same distribution with $\mathbb{E}(y_{t}^{(i)})=\hat{\mu}_{t}$. Thus, according to the law of large numbers, the statistical average of $\{y_{t}^{(i)}\}_{i=1}^{c}$ converges to the expectation $\mu_{t}$ almost surely as $c\rightarrow\infty$.
@@
$\leq\mathbb{E}[(y_{t}^{(i)})^{2}]$ (11)
$\leq\sup_{i,t}\{(y_{t}^{(i)})^{2}\}=\lambda^{2}.~~(\text{A1})$ (12)
@@
$|x_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq|\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+|x_{t}-\mu_{t}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+b+\sigma.~~(\text{A1})$ (16)
@@
$\sum_{t=1}^{T}|x_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq T\cdot(\frac{\lambda}{c}+b+\sigma)=\frac{T(\lambda+\sqrt{c}(b+\sigma))}{\sqrt{c}}.~~(\text{A3})$ (17)
@@
C.1 Variance Analysis of Total Summed Error
The above proof gives a loose bound as an direct result of the law of large number, which does not depend on any assumption on the temporal dependence. In this part, we consider this factor as a pivot for analysing the effects of different forecasting paradigms, identifying the role of our BDO.
@@ C.1 Variance Analysis of Total Summed Error
For simplicity, let the prediction error at each timestep as $t$ $e_{t}:=\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}$, and temporarily omit the effect of the predictive bias $b$ of NFM in A3, i.e., $\mu=\hat{\mu}$. Then we can compute the covariance (autocorrelation) of the errors at any two distinct timesteps $k$ and $h$,
@@
The last inequality holds by assumption A2, which isolates the correlations across different candidates.
@@
Case 1: Temporal Independence (The DO Paradigm): The Direct Output (DO) paradigm is designed to satisfy this assumption of temporal independence, as evidenced in Sec. 2. This implies that the error covariance terms in Eq. 23 are zero. Consequently, the total error variance is simply the sum of the per-timestep variances, While this structure effectively prevents the compounding of errors, it often comes at the cost of the model being ”unaware” of the sequential dynamics within the future horizon, potentially leading to higher per-timestep variance. Furthermore, the variance of the total summed error can only benefit from the combination factor $c$ in reducing the individual uncertainty at each future timestep $\sum_{t=1}^{T}var(e_{t})$.
@@
Case 3: A Synthesis (The BDO Paradigm).: Our Boosted Direct Output paradigm operates as a synthesis of these two extremes. It intentionally violates strict temporal independence by recursively feeding sub-forecasts back into the model, thereby making it aware of causal dependencies within the forecast horizon. However, unlike a pure AR model, it would learn a mapping that minimizes the error covariance. Through hierarchical supervision and its patch-wise output structure inherited from DO, the model is trained to control the accumulation of errors. The goal is to leverage the benefits of modeling temporal structure while constraining the error correlation, effectively learning to make the covariance term in Eq. 23 as small as possible.
@@
Appendix D Implicit structural Loss
@@
An interesting property of the BDO paradigm is the implicit structural loss it induces. This connects our work to recent research on explicit loss engineering, such as positional weighting (xue2023card) and patch-wise structural losses (patchloss). We posit that the BDO learning objective, formed by the weighted sum of losses from hierarchical sub-forecasts, inherently functions as a complex structural loss. This seemingly implicit loss can be seen as the generalized version of the above two.
@@
Appendix E Supplementary Result
In addition to the results from hyperparameter searching, we also present a comparison under another common benchmark configuration: a fixed 96-length look-back window with unified hyperparameters for all prediction horizons. It is important to note, however, that this short-input setting inherently degrades the benefits of our BDO framework. Therefore, to isolate and highlight the effect of our proposals under these constraints, we make minimal adjustments to our best-performing model configuration. We retain most hyperparameters from our main experiments, adjusting only essential parameters like the number of layers for the 720-predict. Under this setup, we compare ReNF against the most recent state-of-the-art models, including OLinear, TQNet, TimeBridge, and CARD.
Models
Re-Bound ReNF TimeBridge OLinear TQNet CARD
- ours
((liu2024timebridge))
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96 0.119 0.161 0.159 0.198 0.171 0.207 0.153 0.190 0.157 0.200 0.150 0.188
192 0.151 0.197 0.206 0.243 0.220 0.250 0.200 0.235 0.206 0.245 0.202 0.238
336 0.188 0.227 0.261 0.285 0.274 0.290 0.258 0.280 0.262 0.287 0.260 0.282
720 0.219 0.249 0.341 0.337 0.352 0.341 0.337 0.333 0.344 0.342 0.343 0.353
Avg. 0.165 0.203 0.242 0.266 0.254 0.272 0.237 0.260 0.242 0.269 0.239 0.265
Electricity
96 0.087 0.173 0.130 0.220 0.139 0.232 0.131 0.221 0.134 0.229 0.141 0.233
192 0.098 0.186 0.146 0.235 0.158 0.250 0.150 0.238 0.154 0.247 0.160 0.250
336 0.106 0.197 0.158 0.249 0.179 0.270 0.165 0.254 0.169 0.264 0.173 0.263
720 0.117 0.209 0.182 0.274 0.208 0.296 0.191 0.279 0.201 0.294 0.197 0.284
Avg. 0.102 0.191 0.154 0.245 0.171 0.262 0.159 0.248 0.165 0.259 0.168 0.258
Traffic
96 0.280 0.176 0.375 0.238 0.386 0.248 0.398 0.226 0.413 0.261 0.419 0.269
192 0.287 0.184 0.398 0.250 0.404 0.257 0.439 0.241 0.432 0.271 0.443 0.276
336 0.298 0.191 0.412 0.257 0.415 0.264 0.464 0.250 0.450 0.277 0.460 0.283
720 0.330 0.208 0.444 0.277 0.456 0.286 0.502 0.270 0.486 0.295 0.490 0.299
Avg. 0.299 0.190 0.407 0.256 0.415 0.264 0.451 0.247 0.445 0.276 0.453 0.282
Solar
96 0.090 0.139 0.173 0.198 0.190 0.208 0.180 0.191 0.173 0.233 0.197 0.211
192 0.106 0.155 0.206 0.225 0.223 0.231 0.209 0.213 0.199 0.257 0.234 0.234
336 0.119 0.170 0.224 0.246 0.242 0.250 0.231 0.229 0.211 0.263 0.256 0.250
720 0.120 0.174 0.231 0.256 0.249 0.256 0.241 0.236 0.209 0.270 0.260 0.254
Avg. 0.109 0.160 0.209 0.231 0.226 0.236 0.215 0.217 0.198 0.256 0.237 0.237
ETTm1
96 0.215 0.273 0.298 0.332 0.313 0.345 0.302 0.334 0.311 0.353 0.316 0.347
192 0.252 0.295 0.350 0.361 0.367 0.378 0.357 0.363 0.356 0.378 0.363 0.370
336 0.277 0.313 0.380 0.385 0.396 0.399 0.387 0.385 0.390 0.401 0.392 0.390
720 0.321 0.340 0.436 0.425 0.457 0.440 0.452 0.426 0.452 0.440 0.458 0.425
Avg. 0.266 0.305 0.366 0.376 0.383 0.391 0.373 0.377 0.377 0.393 0.382 0.383
ETTm2
96 0.138 0.219 0.168 0.246 0.174 0.251 0.169 0.249 0.173 0.256 0.169 0.248
192 0.185 0.254 0.232 0.290 0.242 0.297 0.232 0.290 0.238 0.298 0.234 0.292
336 0.230 0.285 0.289 0.328 0.302 0.338 0.291 0.328 0.301 0.340 0.294 0.339
720 0.299 0.328 0.384 0.385 0.408 0.398 0.389 0.387 0.397 0.396 0.390 0.388
Avg. 0.213 0.272 0.268 0.312 0.282 0.321 0.269 0.314 0.277 0.323 0.272 0.317
ETTh1
96 0.253 0.305 0.360 0.383 0.377 0.392 0.360 0.382 0.371 0.393 0.383 0.391
192 0.288 0.328 0.410 0.414 0.429 0.421 0.416 0.414 0.428 0.426 0.435 0.420
336 0.321 0.347 0.448 0.435 0.470 0.443 0.457 0.438 0.476 0.446 0.479 0.442
720 0.292 0.342 0.442 0.454 0.477 0.473 0.463 0.462 0.487 0.470 0.471 0.461
Avg. 0.289 0.331 0.415 0.422 0.438 0.432 0.424 0.424 0.441 0.434 0.442 0.429
ETTh2
96 0.227 0.286 0.281 0.331 0.287 0.334 0.284 0.329 0.295 0.343 0.282 0.331
192 0.291 0.327 0.359 0.379 0.362 0.383 0.360 0.379 0.367 0.393 0.363 0.381
336 0.323 0.354 0.398 0.412 0.417 0.425 0.409 0.415 0.417 0.427 0.411 0.418
720 0.297 0.349 0.407 0.429 0.436 0.448 0.415 0.431 0.433 0.446 0.416 0.431
Avg. 0.285 0.329 0.361 0.388 0.376 0.398 0.367 0.389 0.378 0.402 0.368 0.390
Table 6: Full results of long-term forecasting of unified hyperparameter. The Re-Bound column denotes the empirical bound discussed in Sec.3.3. The look-back window is fixed to 96 for all baselines. All results are averaged across four different prediction lengths. The best and second-best results are highlighted in red and blue, respectively.
As shown in Table 6, ReNF remains highly competitive against the most recent SOTA models. However, its performance advantage is less pronounced in this setting, particularly in terms of MAE. We hypothesize that when restricted to shorter look-back windows, the initial short-term forecasts in the BDO process are less accurate. Consequently, concatenating these noisier predictions can introduce disruptions that limit the full benefit of our paradigm. From another perspective, while our proposed methods are universally applicable, these results also highlight the value of more specialized techniques in further enhancing the capabilities of Neural Forecasters.
Models
Re-Bound ReNF TimeBridge DUET TimeDistill Timer-XL iTransformer TimeMixer PatchTST Crossformer DLinear
@@ Appendix E Supplementary Result
Table 7: Full results of long-term forecasting of hyperparameter searching. The Re-Bound column denotes the empirical bound discussed in Sec.3.3. The look-back window is searched from $\{336,512,720\}$ for the best performance. Timer-XL uses a 672-length window as the original paper. All results are averaged across four different prediction lengths: $\{96,192,336,720\}$. The best and second-best results are highlighted in red and blue, respectively.
E.1 More Ablations

### v5

@@ B.1 Proof of the Hypothesis
First, we derive the expected bound for each element $y_{t}^{(i)}$ in a candidate forecast $\hat{Y}_{f}^{(i)}=\{y_{1}^{(i)},y_{2}^{(i)},\cdots,y_{c}^{(i)}\}$, by assuming that each candidate forecast at a fixed time $t$, i.e., $\{y_{t}^{(i)}\}_{i=1}^{c}$, are independent samples from the same distribution with $\mathbb{E}(y_{t})=\hat{\mu}_{t}$.
@@
$\leq\mathbb{E}[(y_{t}^{(i)})^{2}]\leq\lambda^{2}.$ (15)
where $\lambda$ represents the upper bound of the second moment $\mathbb{E}[(y_{t}^{(i)})^{2}]$.
@@
$|x_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq|\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+|x_{t}-\mu_{t}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+b+\sigma.$ (19)
@@
$\sum_{t=1}^{N}|x_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq N\cdot(\frac{\lambda}{\sqrt{c}}+b+\sigma)=\frac{N(\lambda+\sqrt{c}(b+\sigma))}{\sqrt{c}}.$ (20)
@@
B.2 Variance Analysis of Total Summed Error
The above proof gives an intuitive bound, which does not depend on any assumption on the temporal dependence. In this part, we consider this factor as a pivot for analysing the effects of different forecasting paradigms, identifying the role of our BDO.
@@ B.2 Variance Analysis of Total Summed Error
For simplicity, let the prediction error at each timestep be $t$ $e_{t}:=\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}$, and temporarily omit the effect of the predictive bias $b$ of NFM, i.e., $\mu=\hat{\mu}$. Then we can compute the covariance (autocorrelation) of the errors at any two distinct timesteps $k$ and $h$,
@@
The last inequality holds by isolating the correlations across different candidates.
@@
Case 1: Temporal Independence (The DO Paradigm): The DO paradigm is designed to (approximately) satisfy the assumption of temporal independence, as evidenced in Sec. 2. This implies that the error covariance terms in Eq. 26 are (nearly) zero. Consequently, the total error variance is simply the sum of the per-timestep variances, While this structure effectively prevents the compounding of errors, it often comes at the cost of the model being ”unaware” of the sequential dynamics within the future horizon, potentially leading to higher per-timestep variance.
@@
Case 3: A Synthesis (The BDO Paradigm).: Our Boosted Direct Output paradigm operates as a synthesis of these two extremes. It intentionally violates strict temporal independence by recursively feeding sub-forecasts back into the model, thereby making it aware of causal dependencies within the forecast horizon. However, unlike a pure AR model, through hierarchical supervision and its patch-wise output structure inherited from DO, the model is trained to control the accumulation of errors. The goal is to leverage the benefits of modeling temporal structure while constraining the error correlation, effectively learning to make the covariance term in Eq. 26 as small as possible.
@@
Appendix C Implicit structural Loss
@@
An interesting property of the BDO paradigm is the implicit structural loss it induces. This connects our work to recent research on explicit loss engineering, such as positional weighting (card) and patch-wise structural losses (patchloss). We posit that the BDO learning objective, formed by the weighted sum of losses from hierarchical sub-forecasts, inherently functions as a complex structural loss. This seemingly implicit loss can be seen as the generalized version of the above two.
@@
Appendix D Full Empirical Results
D.1 Long-term Forecasting Results
We present the full version of Table.2 in Table.6 to show the capability of ReNF in Long-Term time series forecasting.
Models Re-Bound ReNF TimeBridge DUET TimeDistill Timer-XL iTransformer TimeMixer PatchTST Crossformer DLinear
@@ D.1 Long-term Forecasting Results
Table 6: Full results of long-term forecasting of hyperparameter searching. The Re-Bound column denotes the empirical bound discussed in Sec.3.3. The look-back window is searched from $\{336,512,720\}$ for the best performance. Timer-XL uses a 672-length window as in the original paper. All results are averaged across four different prediction lengths: $\{96,192,336,720\}$. The best and second-best results are highlighted.
D.2 Short-term Forecasting Results
To further demonstrate the generality of ReNF, we test it at six supplementary datasets from a recent short-term time series forecasting benchmark (yue2025olinear). The descriptions of these datasets can be found in the Table.5.
Model
ReNF
OLinear
yue2025olinear
TimeMix.
wang2024timemixer
FilterNet
filternet
FITS
fits
DLinear
dlinear
TimeMix.++
timemixer++
Leddam
leddam
CARD
card
Fredformer
fredformer
iTrans.
liu2023itransformer
PatchTST
patchtst
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
CarSales 24 0.263 0.283 0.320 0.302 0.320 0.318 0.318 0.319 0.359 0.347 0.354 0.350 0.323 0.320 0.325 0.322 0.337 0.321 0.319 0.326 0.303 0.312 0.319 0.319
36 0.280 0.298 0.334 0.315 0.332 0.331 0.331 0.330 0.373 0.360 0.368 0.365 0.351 0.348 0.337 0.333 0.348 0.333 0.333 0.335 0.318 0.323 0.332 0.330
48 0.299 0.315 0.347 0.327 0.345 0.343 0.342 0.341 0.385 0.370 0.382 0.379 0.351 0.342 0.351 0.346 0.362 0.345 0.349 0.344 0.331 0.332 0.347 0.344
60 0.315 0.328 0.358 0.337 0.355 0.351 0.352 0.349 0.399 0.385 0.388 0.380 0.363 0.352 0.361 0.353 0.372 0.353 0.359 0.349 0.344 0.342 0.355 0.348
Avg 0.289 0.306 0.340 0.320 0.338 0.336 0.336 0.335 0.379 0.365 0.373 0.368 0.347 0.340 0.343 0.338 0.355 0.338 0.340 0.338 0.324 0.327 0.338 0.335
Power 24 1.268 0.855 1.343 0.870 1.341 0.881 1.410 0.916 1.491 0.944 1.390 0.916 1.340 0.877 1.397 0.909 1.406 0.886 1.410 0.913 1.462 0.924 1.468 0.935
36 1.336 0.880 1.445 0.903 1.420 0.914 1.590 0.968 1.621 0.994 1.518 0.957 1.446 0.920 1.509 0.951 1.506 0.921 1.538 0.953 1.582 0.964 1.593 0.972
48 1.354 0.893 1.559 0.946 1.567 0.963 1.680 1.009 1.775 1.052 1.610 0.995 1.467 0.933 1.646 0.999 1.583 0.957 1.652 1.008 1.696 1.011 1.710 1.020
60 1.383 0.913 1.602 0.971 1.609 0.988 1.776 1.053 1.958 1.122 1.679 1.020 1.626 1.006 1.727 1.043 1.693 1.003 1.752 1.049 1.796 1.061 1.829 1.064
Avg 1.335 0.885 1.487 0.922 1.484 0.937 1.614 0.986 1.711 1.028 1.549 0.972 1.470 0.934 1.570 0.975 1.547 0.942 1.588 0.981 1.634 0.990 1.650 0.998
METR-LA 24 0.613 0.348 0.650 0.337 0.671 0.413 0.670 0.402 0.698 0.416 0.645 0.458 0.617 0.394 0.680 0.405 0.700 0.378 0.676 0.408 0.700 0.413 0.679 0.410
36 0.748 0.398 0.800 0.388 0.841 0.480 0.824 0.471 0.874 0.490 0.785 0.533 0.781 0.457 0.841 0.471 0.874 0.448 0.852 0.477 0.867 0.480 0.845 0.484
48 0.860 0.438 0.905 0.427 0.964 0.531 0.955 0.521 1.013 0.546 0.885 0.585 0.842 0.520 0.963 0.528 1.017 0.498 0.982 0.526 1.017 0.539 0.972 0.536
60 0.950 0.471 0.999 0.457 1.047 0.573 1.050 0.563 1.122 0.589 0.959 0.623 0.958 0.551 1.029 0.556 1.126 0.541 1.084 0.569 1.079 0.572 1.077 0.578
Avg 0.793 0.414 0.838 0.402 0.881 0.499 0.875 0.489 0.927 0.510 0.819 0.550 0.799 0.480 0.878 0.490 0.929 0.466 0.898 0.495 0.916 0.501 0.893 0.502
Website 24 0.155 0.286 0.186 0.306 0.229 0.335 0.273 0.357 0.431 0.469 0.315 0.393 0.231 0.349 0.240 0.345 0.325 0.370 0.216 0.335 0.181 0.305 0.245 0.350
36 0.224 0.343 0.272 0.356 0.361 0.420 0.401 0.441 0.554 0.552 0.385 0.447 0.328 0.397 0.327 0.405 0.428 0.442 0.331 0.411 0.226 0.343 0.370 0.429
48 0.283 0.386 0.365 0.391 0.501 0.507 0.530 0.522 0.694 0.647 0.436 0.486 0.450 0.473 0.446 0.475 0.457 0.478 0.483 0.496 0.263 0.370 0.504 0.513
60 0.267 0.374 0.486 0.481 0.571 0.562 0.630 0.592 0.736 0.673 0.468 0.510 0.525 0.517 0.561 0.549 0.596 0.566 0.556 0.547 0.323 0.410 0.587 0.565
Avg 0.233 0.348 0.327 0.383 0.415 0.456 0.458 0.478 0.604 0.585 0.401 0.459 0.384 0.434 0.393 0.443 0.451 0.464 0.396 0.447 0.248 0.357 0.426 0.464
SP500 24 0.150 0.270 0.155 0.271 0.159 0.288 0.181 0.317 0.193 0.334 0.189 0.330 0.172 0.305 0.175 0.308 0.156 0.276 0.181 0.315 0.180 0.309 0.164 0.298
36 0.204 0.322 0.209 0.317 0.218 0.343 0.224 0.341 0.259 0.389 0.250 0.363 0.227 0.344 0.232 0.358 0.206 0.319 0.239 0.365 0.225 0.346 0.221 0.341
48 0.251 0.355 0.258 0.358 0.264 0.367 0.280 0.384 0.324 0.439 0.291 0.398 0.272 0.383 0.276 0.388 0.258 0.354 0.283 0.394 0.275 0.383 0.278 0.397
60 0.295 0.386 0.305 0.387 0.322 0.416 0.332 0.416 0.391 0.486 0.377 0.475 0.319 0.413 0.325 0.423 0.303 0.385 0.341 0.438 0.322 0.418 0.321 0.409
Avg 0.225 0.333 0.231 0.333 0.241 0.353 0.254 0.365 0.291 0.412 0.277 0.391 0.247 0.361 0.252 0.369 0.231 0.333 0.261 0.378 0.250 0.364 0.246 0.361
NASDAQ 24 0.114 0.211 0.121 0.216 0.122 0.221 0.130 0.230 0.140 0.244 0.155 0.274 0.132 0.233 0.125 0.222 0.124 0.220 0.128 0.226 0.137 0.237 0.127 0.224
36 0.155 0.253 0.163 0.261 0.183 0.279 0.175 0.273 0.184 0.284 0.196 0.306 0.177 0.278 0.174 0.271 0.167 0.266 0.170 0.268 0.184 0.280 0.174 0.269
48 0.196 0.290 0.205 0.296 0.200 0.298 0.224 0.314 0.234 0.324 0.244 0.344 0.216 0.311 0.222 0.312 0.218 0.307 0.218 0.306 0.229 0.318 0.225 0.314
60 0.236 0.321 0.259 0.336 0.238 0.328 0.259 0.340 0.282 0.357 0.318 0.401 0.249 0.337 0.264 0.341 0.264 0.341 0.262 0.339 0.279 0.352 0.265 0.339
Avg 0.175 0.269 0.187 0.277 0.186 0.281 0.197 0.289 0.210 0.302 0.228 0.331 0.193 0.290 0.196 0.286 0.193 0.284 0.194 0.285 0.207 0.297 0.198 0.286
Table 7: Full results for the short-term forecasting. We use the look-back window with length $T=36$ to predict lengths $\left\{24,36,48,60\right\}$. The best results and second-best results are highlighted.
As shown in Table 7, ReNF remains highly competitive against recent state-of-the-art models in short-term forecasting tasks. However, the performance advantage is less pronounced than in the LTSF setting, a result that aligns with the design principles of BDO. First, when restricted to the shorter look-back windows typical of short-term tasks, the initial forecasts in the BDO process are inherently less accurate. Consequently, recursively combining these noisier predictions can offset the benefits of the paradigm. Second, the hierarchical curriculum learning of BDO yields diminishing returns on short horizons; the strategy of decomposing a short prediction interval into even smaller sub-forecasts becomes redundant. We provide a more rigorous verification of these dynamics in Sec. E.1, where we conduct ablation studies on the BDO module using the M4 and PEMS datasets.
From another perspective, while our proposed methods are universally applicable, these results also highlight the value of more specialized or refined extensions of our proposals in further enhancing the capabilities of BDO in various scenarios.
Appendix E Further Ablations